OpenStudy (anonymous):

What is 40% of 3,000?

mathslover (mathslover):

can u solve this? : \[\large{\frac{40}{100}*3000}\]

OpenStudy (anonymous):

uh i dont understand multiplying fractions

mathslover (mathslover):

hmn : \[\large{\frac{a}{b}*c=\frac{a*c}{b}}\]

mathslover (mathslover):

ram is helping u best of luck

OpenStudy (anonymous):

ok

mathslover (mathslover):

\[\large{\frac{4*10}{1*100}*3*1000}\] \[\large{\frac{4}{10}*3000}\] \[\large{4*300=1200}\]